### v1

@@ Abstract
Recent research in active learning, and more precisely in uncertainty sampling, has focused on the decomposition of model uncertainty into reducible and irreducible uncertainties. In this paper, we propose to simplify the computational phase and remove the dependence on observations, but more importantly to take into account the uncertainty already present in the labels, i.e. the uncertainty of the oracles. Two strategies are proposed, sampling by Klir uncertainty, which addresses the exploration-exploitation problem, and sampling by evidential epistemic uncertainty, which extends the reducible uncertainty to the evidential framework, both using the theory of belief functions.
@@
Active Learning Uncertainty sampling Belief Functions
@@ 1 Introduction
For reasons of efficiency, cost or energy reduction in machine learning or deep learning, one of the important issues is related to the amount of data and in some cases, to the amount of labelled data. Active learning [19] is a part of machine learning in which the learner can choose which observation to label in order to work with only a fraction of the labeled dataset to reduce the labeling cost. For this purpose, the learner uses a strategy that allows it to select only certain observations that will then be labeled. Among all the proposed strategies in the literature [19, 1] one of the best known is sampling by uncertainty [15].
In uncertainty sampling, the learner selects the instances for which it is most uncertain. The measures used to quantify this uncertainty, such as entropy, are up to now probabilistic. In this paper, we propose to use a broader framework of uncertainty that generalizes probabilities.
As proposed in recent papers [10, 11, 18] the uncertainty can be decomposed into two interesting terms: the epistemic and the aleatoric uncertainties. Aleatoric uncertainty arises from the stochastic property of the event and is therefore not reducible, whereas epistemic uncertainty is related to a lack of knowledge and can be reduced. Proposed calculations depend on the model prediction but also on the observations. We suggest in this paper, to get rid of the direct dependence on the observations and to use only the model output for similar results. This representation also addresses the exploration-exploitation problem in active learning, with the possibility of choosing one or the other, or even a compromise as in [2].
The labeling process is often carried out by humans [17, 7]; without making any difference between a label given by someone who has hesitated for a long time and a label given by someone who has no doubt, and therefore uncertainty may already exist in the labels. This information is not taken into account in most models and sampling strategies. In the case of supervised classification, several models are now able to handle these uncertain labels [5, 6, 4, 23]. The main objective, in addition to not being dependent on observations and to address the problem of exploration-exploitation, is to take into account in the sampling, the uncertainty already present in the labels.
Given the above, we propose in this paper two uncertainty sampling strategies capable of representing a decomposition of the model uncertainties with regard to the uncertainty already present in the labels.
The first strategy is based upon two different uncertainties, the discord (how self-conflicting the information is) and non-specificity (how ignorant the information is) in the model output. The second strategy extends the epistemic uncertainty to the evidential framework and to several classes, thus simplifying the computation.
@@
The paper is organized as follows; section 2 introduces some important notions of imperfect labeling and the modeling of these richer labels using the theory of belief functions. The usual uncertainty sampling approach [15] is also recalled and section 3 describes the separation between aleatoric and epistemic uncertainties. Section 4 presents the two new proposed strategies, section 5 shows an application on a real world dataset, then section 6 discusses and concludes the article. The experiments performed in this paper are described in supplementary materials, to avoid lengthy explanations, since the purpose of the paper does not lie in this part. Furthermore, uncertainties are mapped on 2D representations but the objective is to later serve active learning.
@@ 2 Preliminaries
In this section, we introduce some general knowledge useful to understand the rest of the paper, starting with rich labels, modeled by the theory of belief functions and ending with the classical approach of sampling by uncertainty.
Imperfect labeling -
@@
Theory of belief functions -
The theory of belief functions [3, 20], is used in this study to model uncertainty and imprecision for labeling and prediction. Let $\Omega=\{\omega_{1},\ldots,\omega_{M}\}$ be the frame of discernment for $M$ exclusive and exhaustive hypotheses. It is assumed that only one element of $\Omega$ is true (closed-world assumption) [21]. The power set $2^{\Omega}$ is the set of all subsets of $\Omega$. A mass function assigns the belief that a source may have about the elements of the power set of $\Omega$, such that the sum of all masses is equal to $1$.
@@
On decision level, the pignistic probability $BetP$ [21] helps decision making on singletons:
@@
There are other possible combinations that are more common than the mean, many of which are listed in [14].
$\bullet$ Example 1:
@@
$\bullet$ Example 2: An observation labeled “Cat or Dog” by a source can be modeled by the mass function $m_{2}$ such as: $m_{2}(\{Cat,Dog\})=1$ and $m_{2}(A)=0$, $\forall A\in 2^{\Omega}\backslash\{Cat,Dog\}$.
$\bullet$ Example 3: The average mass function $\bar{m}$ of $m_{1}$ and $m_{2}$ is: $\bar{m}(\{Cat\})=0.5$, $\bar{m}(\{Cat,Dog\})=0.5$ and $\bar{m}(A)=0$ for all other subsets $A$ in $2^{\Omega}$. Its pignistic probability $BetP$, used for decision making is: $BetP(\{Cat\})=0.75$ and $BetP(\{Dog\})=0.25$.
Uncertainty sampling -
Active learning iteratively builds a training set by selecting the best instances to label. The principle is, for a given performance or a given budget, to label as few observations as possible. Among all the strategies proposed in the literature [19] one of the best known methods is uncertainty sampling [13], where the function that defines the instances to be labeled maximizes the uncertainty related to the model prediction as described below.
@@
with $p(\omega|x)$ the probability for $x$ to belong to the class $\omega$, given by the model. Other uncertainty criteria exist, it is common to use the least confidence measure:
@@
Measuring the uncertainty of a model to predict the class of some observations can be useful to find the areas of uncertainty in a space.
@@
Figure 1: Three 2-class datasets with areas of model uncertainty.
Given the model and one of the uncertainty criteria, we can compute the uncertainty of any point in space. For each dataset, the areas of uncertainty of the model are represented, with more red for more uncertainty. It is remarkable that these uncertainty areas can be compared to the decision boundaries of the model. Often, the closer the observation is to the decision boundary, the less confident the model is about its prediction.
Uncertainty sampling consists of choosing the observation for which the model is least certain of its prediction. This is one of the basis of active learning, however, other methods allow to extract more information about this uncertainty which leads to the decomposition into epistemic and aleatoric uncertainties.
@@ 3 On the interest and limits of epistemic and aleatoric uncertainties for active learning
The uncertainty $\mathcal{U}(x)$ can be separated into two uncertainties [9], one reducible and the other irreducible. The example [^1] This example is taken from Eyke Hüllermeier’s talk “Representation and Quantification of Uncertainty in Machine Learning” at the LFA2022 conference. In our example the word tails is written in Finnish, the word heads is called “Kruuna”. of Figure 2 shows these two types of uncertainties, on 1(a) the result of a coin toss is uncertain and it is not possible to generate more knowledge to predict that the coin will flip heads or tails, this ignorance is called aleatoric uncertainty. On 1(b) either heads or tails is written in Finnish, it is an uncertainty that can be resolved by learning this language, it is called epistemic uncertainty.
@@
Figure 2: Representation of aleatoric and epistemic uncertainties through the tossing of a coin and the word “heads” or “tails” written in Finnish.
Being able to model these two uncertainties can help delimit where it is more interesting to provide knowledge and where it is useless. The total uncertainty $\mathcal{U}(x)$ is often represented as the sum of the epistemic uncertainty $\mathcal{U}_{e}(x)$ and the aleatoric uncertainty $\mathcal{U}_{a}(x)$: $\mathcal{U}(x)=\mathcal{U}_{e}(x)+\mathcal{U}_{a}(x)$.
For a two-class problem $\Omega=\{0,1\}$, it is proposed in [18] to model this uncertainty (here under the [15] formalism) by computing the plausibility $\pi$ of belonging to each of the two classes with the following formula, according to a probabilistic model $\theta$:
@@
[…]
The epistemic uncertainty is then high when the two classes are very plausible while the aleatoric uncertainty is high when the two classes are implausible:
@@
Figure 3: Sample with areas of uncertainty according to the uncertainty sampling and three possible datasets based on the observations available in 2(a).
@@
Figure 4: Areas of uncertainty on 2(a) for epistemic and aleatoric uncertainties.
An example is shown in Figure 3, a two-class dataset is shown in 2(a) and the areas of model uncertainty are shown in 2(b) according to the uncertainty sampling presented in the previous section. An horizontal line can be distinguished where the model uncertainty is highest. However, the sample represented in 2(a), shows that part of the uncertainty can be removed more easily by adding observations. In the same figure, three different datasets show how the sample can evolve by adding observations. Whatever the final distribution, the uncertainty on the left is not very reducible, while the uncertainty on the right can be modified by adding knowledge.
These two uncertainties can be calculated using equation (8), and are shown in Figure 4. The aleatoric uncertainty, and therefore irreducible, is represented in 3(a) and the epistemic uncertainty, reducible, is represented in 3(b). The total uncertainty is then the sum of the two 3(c). The goal here is to use only the epistemic uncertainty, to know the areas where the model can learn new knowledge and where it will have more impact.
Using epistemic uncertainty as a sampling strategy is not reductive since it provides similar areas of uncertainty to those used previously, where epistemic and aleatoric uncertainty are indistinguishable.
Such information can be useful to find areas of reducible uncertainty, but it is not compatible with richer labels that also contain uncertainty. The way to compute this epistemic uncertainty is also dependent on the observations in addition to the model (i.e. the method could be oversimplified as: the model defines its zones of uncertainty, in which we look for the location with the smallest number of observations to define the reducible uncertainty.). Furthermore, the exploration-exploitation problem is not fully addressed. This leads to the next section in which two uncertainty sampling strategies for rich labels are proposed, they are also extended to several classes.
@@ 4 Richer labels and multiple classes
In this section, we propose two uncertainty sampling strategies, with a simplified calculation phase, able to deal with richer labels and no longer directly dependent on the observations but only on the model prediction [^2] The uncertainty is no longer directly dependent on the observations, but the model still is.. We also propose a natural extension for a number of classes higher than two. The first method uses discord and non-specificity to map uncertainty in order to address the exploration-exploitation problem. The second method extends the epistemic and aleatoric uncertainties to rich labels, also simplifying the computation phase.
From there, a label can be uncertain and imprecise, which means that additional information on ignorance is represented. Figure 5 shows how these labels are represented in this document, the darker the dot, the less ignorance the label contains (e.g. I’m sure this is a dog), the lighter the dot, the more ignorance it contains (e.g. I have no idea between dog and cat).
Figure 5: Observations on two dimensions with their rich labels, the darker the point, the more certain and precise its label.
@@ 4.1 Discord and non-specificity: Klir uncertainty
In the framework of belief functions, discord and non-specificity are tools that allow to model uncertainty, we propose to use Klir’s representation [12] for uncertainty sampling, some bridges can be made with epistemic and aleatoric uncertainty.
4.1.1 Discord
is here applied to the output of a model capable of making an uncertain and imprecise prediction [^3] The Evidential $K$-nearest Neighbors model [5] is considered to illustrate the examples, which may vary depending on the model used.. It represents the amount of conflicting information in the model’s prediction and is calculated with the following formula:
@@ 4.1.1 Discord
with $m$ a mass function, or the output of the model (see section 2).
Figure 6 represents three different cases where the discord varies, from high discordance where labels around the central point (the observation to label) highly disagree 5(a), to low discordance where each of the labels is in agreement 5(c).
4.1.2 Non-Specificity
allows to quantify the degree of ignorance of the model, the higher it is, the more imprecise the response of the model, it is calculated with:
$N(m)=\sum_{A\subseteq\Omega}m(A)\text{ log}_{2}(|A|).$ (10)
The same Figure 6 also represents three different cases of non-specificity, in 5(d) the non-specificity is low as there are relevant sources of information next to the observation to be labelled, in 5(e) the non-specificity increases the further away the elements are from the observation and in 5(f) the non-specificity is also high because the nearby sources of information are themselves ignorant.
@@ 4.1.2 Non-Specificity
(d) Low non-specificity
(e) High non-specificity
(f) High non-specificity
Figure 6: Three degrees of discord and three degrees of non-specificity in the center.
4.1.3 Klir uncertainty
is then derived from discord and non-specificity, it is used here for uncertainty sampling by adding the two previous formulas:
@@ 4.1.3 Klir uncertainty
with $N(x)$ and $D(x)$ respectively the non-specificity and discord of the model in $x$. Klir [12] proposes to use the same weight for discord and non-specificity, but in [4] a parameter $\lambda\in[0,1]$ is introduced and allows to bring more weight to non-specificity (we propose to use it for more exploration) or to discord (for more exploitation):
@@
This formula has the advantage of identifying the total uncertainty as well as the reducible one, but also of taking into account the uncertainty already present in the labels and of being adjustable for more exploration or exploitation. Figure 7 shows a dataset with two areas of uncertainty 6(a), on the right an area with a lack of data and on the left an area where labels are more ignorant. The uncertainty sampling, using Shannon’s entropy (4) or the least confidence measure (5) is not able to see either of these two areas 6(b). The epistemic uncertainty (8) is able to distinguish the uncertainty related to the arrangement of the observations in space (i.e. the uncertainty on the right) but not the uncertainty related to the ignorance of the sources 6(c).
@@
Figure 7: An imperfectly labeled dataset 6(a) with the areas of uncertainty according to uncertainty sampling and epistemic uncertainty.
(a) Non-specificity
(b) Discord
(c) Klir
Figure 8: Areas of uncertainty corresponding to the dataset 6(a) according to the non-specificity, the discord and the total uncertainty defined by Klir.
(a) More exploitation
(b) Compromise
(c) More exploration
Figure 9: Areas of Klir uncertainty, modifying the amount of non-specificity and discord. With $\lambda=0.1$, more discord is taken into account, with $\lambda=0.5$, discord and non-specificity are used as much and with $\lambda=0.9$, more non-specificity is taken into account.
The proposal of using Klir uncertainty for sampling (discord and non-specificity) allows to represent each of these uncertainties. Figure 8 shows the areas of non-specificity 7(a), of discord 7(b) and Klir uncertainty 7(c).
Klir uncertainty can then be used for uncertainty sampling in active learning, it is also possible to vary the result for more exploration or more exploitation by modifying $\lambda$. Figure 9 shows the areas of uncertainty for different values of $\lambda$, more discord on the left to more non-specificity on the right.
We have proposed here to use Klir’s uncertainty in sampling, which allows to represent some unknown uncertainties areas in active learning related to rich labels. The method is no longer dependent on the observations, but only on the prediction of the model and the exploration-exploitation problem is addressed thanks to the $\lambda$ parameter. Even though discord may recall aleatoric uncertainty (non-reducible) and non-specificity may recall epistemic uncertainty (reducible). These notions are not quite equivalent. Therefore, in the following section we also propose an extension of epistemic (and aleatoric) uncertainty for rich labels and for several classes.

### v2

@@ Abstract
Recent studies in active learning, particularly in uncertainty sampling, have focused on the decomposition of model uncertainty into reducible and irreducible uncertainties. In this paper, the aim is to simplify the computational process while eliminating the dependence on observations. Crucially, the inherent uncertainty in the labels is considered, i.e. the uncertainty of the oracles. Two strategies are proposed, sampling by Klir uncertainty, which tackles the exploration-exploitation dilemma, and sampling by evidential epistemic uncertainty, which extends the concept of reducible uncertainty within the evidential framework, both using the theory of belief functions. Experimental results in active learning demonstrate that our proposed method can outperform uncertainty sampling.
@@
Active Learning, Uncertainty sampling, Belief Functions
@@ 1 Introduction
Active Learning (AL) - For reasons of efficiency, cost or energy reduction in machine learning or deep learning, one of the important issues is related to the amount of data and in some cases, to the amount of labeled data. Active learning [32] is a part of machine learning in which the learner can choose which observation to label in order to work with only a fraction of the labeled dataset to reduce the labeling cost. While primarily used for cost reduction [14], active learning finds application in various domains like anomaly detection, as seen in [2] and [26]. The essence of active learning lies in empowering the learner to strategically label selected observations. Among all the proposed strategies in the literature [32, 3], one of the most recognized is uncertainty sampling [25, 28].
Uncertainty Quantification (UQ) - finds applications across various fields, including medical image analysis, as discussed in a review by [19], and in-depth exploration of deep learning applications and techniques [1], such as recent evidential deep learning [31]. Concerning frameworks for uncertainty, numerous methods exist for quantifying uncertainty, with many applied to credal sets as reviewed by [21], or evidential entropies  [8]. Despite being described several years ago [18], recent literature [20, 22, 30, 5] distinguishes two main types of uncertainty: epistemic and aleatoric. Aleatoric uncertainty arises from the stochastic property of the event and is therefore not reducible, whereas epistemic uncertainty is related to a lack of knowledge and can be reduced. Most proposed calculations hinge not only on the model predictions but also on parameter estimations derived directly from the observations themselves.
(AL) $\cup$ (UQ) - What is the issue? - In uncertainty sampling, the learner selects the instances for which it is most uncertain. Until recently, the literature has mostly proposed measures to quantify this uncertainty, such as entropy, in a probabilistic form. But this kind of uncertainty cannot exploit and capture the difference between a label given by someone who has hesitated for a long time and a label given by someone who has no doubt, and therefore uncertainty that may already exist in the labels. In this paper, we propose to use evidential reasoning within the context of active learning. Moreover we propose eliminating direct dependence on the observations and advocating for solely utilizing the model output to achieve a comparable decomposition (i.e epistemic-aleatoric) of uncertainty. This also tackles the exploration-exploitation issue in active learning, with the possibility of choosing one or the other, or even a compromise as suggested by [4]. This paper will use the theory of belief functions that generalises probabilities and will use it in the context of active learning.
Figure 1: Illustrating the exploration-exploitation dilemma in active learning: complete dataset vs. active learning iterations.
Note on Exploration-Exploitation Dilemma - In Figure 1, we present a visual depiction of the exploration-exploitation dilemma. In a 2D classification task, the left panel portrays a scenario where all observations are labeled, while the right panel reflects the outcome after iterative labeling rounds, resulting in sparse observations. Should the sampling strategy heavily favor exploitation, as denoted by the blue line representing the classifier, it is evident that the classifier will neglect further investigation into the top left corner. Consequently, upon encountering the red examples in this unexplored territory later in the process, the classifier’s performance undergoes a significant decline, necessitating extensive parameter adjustments. Conversely, a purely exploratory strategy prolongs the duration required to unveil critical patterns, including the “two red patterns” mentioned. Thus, an optimal strategy must delicately balance exploration and exploitation to navigate this trade-off effectively.
Contributions of this paper - The primary goal of this paper is to consider the uncertainty inherent in the labels (introduced by the entities labeling observations, referred to as ”Oracles” in active learning), and to address the exploration-exploitation dilemma, during sampling. To this end we propose two uncertainty sampling strategies capable of representing a decomposition of the model uncertainties with regard to the uncertainty already present in the labels: (i) a first strategy which is based upon two different uncertainties, discord - how self-conflicting the information is - and non-specificity - how imprecise the information is - in the model output; and (ii) a second strategy which extends the epistemic uncertainty to the evidential framework and to several classes, thus simplifying the computation. To succeed in this challenge we use evidential models able to handle such uncertain labels, such as [10, 13, 9, 36]. By doing this, one can effectively distinguish and account for the difference between a label provided by someone who has hesitated extensively and a label given by someone who has no doubts. As a result, we can identify and quantify the uncertainty inherent in the labels themselves.
@@
Organization - The paper is organized as follows: section 2 introduces some important notions of imperfect labeling and the modeling of these richer labels using the theory of belief functions. The conventional uncertainty sampling approach is also recalled and section 3 describes the separation between aleatoric and epistemic uncertainties. Section 4 introduces the two new proposed strategies and section 5 presents the experiments [^1] For details on experiments conducted in theoretical sections, visit: https://anonymous.4open.science/r/evidential-uncertainty-sampling-D453, first on a real world dataset with rich labels and then in active learning to highlight the relevance and efficacy of the proposed method. Section 6 discusses the encountered limits and section 7 concludes the article.
@@ 2 Preliminaries
In this section, we provide foundational knowledge essential for understanding the rest of the paper, beginning with rich labels, which are characterized by the theory of belief functions, and concluding with the classical method of uncertainty sampling.
2.1 Imperfect labeling
@@
2.2 Theory of belief functions
The theory of belief functions introduced by [6] and [33], is used in this study to model uncertainty and imprecision for labeling and prediction. Let $\Omega=\{\omega_{1},\ldots,\omega_{M}\}$ be the frame of discernment for $M$ exclusive and exhaustive hypotheses. In supervised learning, this refers to the labels (i.e., classes), or the output space. It is assumed that only one element of $\Omega$ is true (closed-world assumption [34]). The power set $2^{\Omega}$ is the set of all subsets of $\Omega$. A mass function assigns the belief that a source may have about the elements of the power set of $\Omega$, such that the sum of all masses is equal to $1$.
@@ 2.2 Theory of belief functions
On decision level, the pignistic probability $BetP$ of [34] helps decision making on singletons:
@@
There are other possible combinations that are more common than the mean, many of which are listed by [27].
Example 1.
@@
Example 2.
An observation labeled “Cat or Dog” by a source can be modeled by the mass function $m_{2}$ such as: $m_{2}(\{Cat,Dog\})=1$ and $m_{2}(A)=0$, $\forall A\in 2^{\Omega}\backslash\{Cat,Dog\}$.
Example 3.
The average mass function $\bar{m}$ of $m_{1}$ and $m_{2}$ is: $\bar{m}(\{Cat\})=0.5$, $\bar{m}(\{Cat,Dog\})=0.5$ and $\bar{m}(A)=0$ for all other subsets $A$ in $2^{\Omega}$. Its pignistic probability $BetP$, used for decision making, is given as follows: $BetP(\{Cat\})=0.75$ and $BetP(\{Dog\})=0.25$.
2.3 Uncertainty sampling
Active learning iteratively builds a training set by selecting the best instances to label. The principle is to label as few observations as possible for a given performance or to achieve the best possible performance within a given budget. Among all the strategies proposed in the literature [32] one of the best known methods is uncertainty sampling [25], where the function that defines the instances to be labeled maximizes the uncertainty related to the model prediction as described below.
@@ 2.3 Uncertainty sampling
with $p(\omega|x)$ the probability for $x$ to belong to the class $\omega$, given by the model. Other common uncertainty criteria include the least confidence measure:
@@
Measuring the uncertainty of a model to predict the class of some observations can be useful to identify the areas of uncertainty in a space.
@@
Figure 2: Visualization of uncertainty areas in two-dimensional datasets.
Given the model and one of the uncertainty criteria [^2] From now on, the model used is $K$-NN ($K$-Nearest Neighbors), with a probabilistic output and on the distance-weighted version available with scikit-learn [29], every other parameters are scikit-learn default parameters. The uncertainty used is the least confidence measure given in equation (5)., we can compute the uncertainty of any point in space. For each dataset, the areas of uncertainty of the model are represented, with more red for more uncertainty. It is remarkable that these uncertainty areas can be compared to the decision boundaries of the model. Often, the closer the observation is to the decision boundary, the less confident the model is about its prediction.
Uncertainty sampling consists of choosing the observation for which the model is the least certain of its prediction. This is one of the basis of active learning, however, other methods allow to extract more information about this uncertainty which leads to the decomposition into epistemic and aleatoric uncertainties.
@@ 3 On the interest and limits of epistemic and aleatoric uncertainties for active learning
The uncertainty $\mathcal{U}(x)$ can be divided into two types, as outlined by [18]: one is reducible, and the other is irreducible. The example provided in Figure 3 illustrates these distinctions. In Fig. 3(a), the outcome of a coin toss is uncertain, and it is impossible to gain further knowledge to predict whether the coin will land heads or tails. This lack of knowledge is referred to as aleatoric uncertainty. On the other hand, in Fig. 3(b), a word in Finnish [^3] In the example, the word “tails” is written in Finnish, the word “heads” is called Kruuna. representing either heads or tails is shown. This uncertainty can be resolved by learning the language, making it epistemic uncertainty.
@@
Figure 3: Illustration of reducible and irreducible uncertainties in a coin toss experiment (and a Finnish word representation).
Being able to model these two uncertainties can help to delimit where it is more interesting to provide knowledge and where it is useless. The total uncertainty $\mathcal{U}(x)$ is often represented as the sum of the epistemic uncertainty $\mathcal{U}_{e}(x)$ and the aleatoric uncertainty $\mathcal{U}_{a}(x)$: $\mathcal{U}(x)=\mathcal{U}_{e}(x)+\mathcal{U}_{a}(x)$.
In a two-class problem where $\Omega=\{0,1\}$, it is suggested by [30] to model this uncertainty, here under the formalism of [28], by computing the plausibility $\pi$ of belonging to each of the two classes with the following formula, based on a probabilistic model $\theta$:
@@
$\pi_{\Theta}(\theta)=\frac{L(\theta)}{L(\hat{\theta})}.$ (7)
The epistemic uncertainty is then high when the two classes are very plausible [^4] The notion of plausibility within the theory of belief functions used in the proposed methods differs from the one presented here and will be discussed in greater detail in section 4. while the aleatoric uncertainty is high when the two classes are implausible:
@@
Figure 4: Visualization of model uncertainty and sample evolution in two-class datasets.
@@
Figure 5: Representation of aleatoric and epistemic uncertainties in model predictions according to Fig. 4(a).
An example is shown in Figure 4, a two-class dataset is shown in Fig. 4(a) and the areas of model uncertainty are shown in Fig. 4(b) according to the uncertainty sampling presented in the previous section.
An horizontal line can be distinguished where the model uncertainty is the highest. However, the sample represented in Fig. 4(a), shows that part of the uncertainty can be removed more easily by adding observations. In the same figure, three different datasets show how the sample can evolve by adding observations. Whatever the final distribution is, the uncertainty on the left is not very reducible, while the uncertainty on the right can be modified by adding knowledge.
These two uncertainties can be calculated using equation (8), and are shown in Figure 5. The aleatoric uncertainty, and therefore irreducible, is represented in Fig. 5(a) and the epistemic uncertainty, reducible, is represented in Fig. 5(b). The total uncertainty is then the sum of the two, Fig. 5(c). Here the goal is to only use the epistemic uncertainty, to know the areas where the model can learn new knowledge and where it will have more impact.
Using epistemic uncertainty as a sampling strategy is not reductive since it theoretically provides similar areas of uncertainty to those used previously when epistemic and aleatoric uncertainties are indistinguishable. But this statement is based on the sum decomposition of total uncertainty, which has recently been questioned.
This information is valuable for identifying areas of reducible uncertainty. However, it is not compatible with richer labels containing uncertainty. Computing this epistemic uncertainty also relies on observations in addition to the model. Essentially, the model defines its zones of uncertainty and seeks locations with the fewest observations to define the reducible uncertainty. Moreover, the exploration-exploitation problem is not fully addressed. This leads to the next section where two uncertainty sampling strategies for rich labels are proposed, extending to multiple classes.
@@ 4 Richer labels and multiple classes
In this section, we propose two uncertainty sampling strategies with a simplified calculation phase, capable of handling richer labels. These strategies are no longer directly dependent on observations but only on the model prediction [^5] The uncertainty no longer depends on observations, but the model does.. We also propose a natural extension for a number of classes higher than two. The first method uses discord and non-specificity to map uncertainty in order to address the exploration-exploitation problem. The second method extends the epistemic and aleatoric uncertainties to rich labels, also simplifying the computation phase.
From there, a label can be uncertain and imprecise, which means that additional information on ignorance is represented. Figure 6 illustrates how labels are represented in this document: the darker the dot, the less ignorance the label contains (e.g., I’m sure this is a dog); the lighter the dot, the more ignorance it contains (e.g., I have no idea between dog and cat). It is important to note that labels are no longer “hard”, but modeled by a belief function, which allows such a representation.
Figure 6: Rich label representation: observations on two dimensions with varying ignorance.
@@ 4.1 Discord and non-specificity: Klir uncertainty
In the framework of belief functions, discord and non-specificity are tools that allow to model uncertainty, we propose to use [23]’s representation for uncertainty sampling, with potential connections to epistemic and aleatoric uncertainty.
Discord
It is here applied to the output of a model capable of making an uncertain and imprecise prediction [^6] From now, the Evidential $K$-nearest Neighbors model of [10] is considered.. Discord represents the amount of conflicting information in the model’s prediction. It is, for most models, at its maximum closest to the decision boundary and is calculated using the following formula:
@@
with $m$ a mass function, or the output of the model (see section 2.2).
@@
(d) Low non-spec.
(e) High non-spec.
(f) High non-spec.
Figure 7: Quantifying discord and non-specificity in model uncertainty at the central point.
Figure 7 illustrates three different cases where discord varies: from high discordance, where labels around the central point (the observation to label) highly disagree in Fig. 7(a), to low discordance, where each label is in agreement in Fig. 7(c).
Non-Specificity
Non-Specificity quantifies the degree of imprecision of the model [12]. This information may be inferred because the model lacks data or because the oracle labeling the instances is itself ignorant. The higher it is, the more imprecise the model’s response, it is calculated with:
$N(m)=\sum_{A\subseteq\Omega}m(A)\text{ log}_{2}(|A|).$ (10)
The same Figure 7 also represents three different cases of non-specificity, in Fig. 7(d) the non-specificity is low as there are relevant sources of information next to the observation to be labeled, in Fig. 7(e) the non-specificity increases the further away the elements are from the observation and in Fig. 7(f) the non-specificity is also high because the nearby sources of information are themselves ignorant.
Klir uncertainty
This uncertainty is derived from discord and non-specificity, it is used here for uncertainty sampling by adding the two previous formulas:
@@
with $N(x)$ and $D(x)$ respectively the non-specificity and discord of the model in $x$. [23] propose to use the same weight for discord and non-specificity, but [9] introduce a parameter $\lambda\in[0,1]$ that allows to bring more weight to non-specificity (we propose to use it for more exploration) or to discord (for more exploitation):
@@
This formula has the advantage of identifying the total uncertainty as well as the reducible one, but also of taking into account the uncertainty already present in the labels and of being adjustable for more exploration or exploitation. Figure 8 depicts a dataset with two areas of uncertainty: on the right, an area with a lack of data, and on the left, an area where labels are more ignorant. The uncertainty sampling, using Shannon’s entropy (4) or the least confidence measure (5) is not able to see either of these two areas, Fig. 8(b). The epistemic uncertainty (8) is able to distinguish the uncertainty related to the arrangement of the observations in space (i.e. the uncertainty on the right) but not the uncertainty related to the ignorance of the sources, Fig. 8(c).
@@
(d) Non-specificity
(e) Discord
(f) Klir
(g) + Exploitation
(h) Compromise
(i) + Exploration
Figure 8: An imperfectly labeled dataset: exploring uncertainty areas through sampling strategies, epistemic uncertainty, and proposed non-specificity, discord, and total Klir uncertainty, alongside the potential for exploration-exploitation compromise.
The proposal of using Klir uncertainty for sampling (discord and non-specificity) allows to represent each of these uncertainties. The second line shows the areas of non-specificity Fig. 8(d), of discord Fig. 8(e) and Klir uncertainty Fig. 8(f).
Klir uncertainty can then be used for uncertainty sampling in active learning, it is also possible to vary the result for more exploration or more exploitation by modifying $\lambda$. The last line shows the areas of uncertainty for different values of $\lambda$, more discord on the left Fig. 8(g) to more non-specificity on the right Fig. 8(i).
We have proposed here to use Klir’s uncertainty in sampling, which allows to represent some uncertainties areas in active learning related to rich labels. The method is no longer dependent on the observations, but only on the prediction of the model and the exploration-exploitation problem is addressed thanks to the $\lambda$ parameter. Even though discord may recall aleatoric uncertainty (non-reducible) and non-specificity may recall epistemic uncertainty (reducible), these notions are not equivalent. Therefore, in the following section we also propose an extension of epistemic (and aleatoric) uncertainty for rich labels and for several classes.